# Forecasting mortality rates with functional signatures

### v1

@@ 2.6.1 Point forecast
Similar to that of the HU model, the time dependent coefficients are extrapolated by fitting them into univariate time series $\{\hat{\beta}_{t,k}\}_{t=1}^{n}$. We preserve the choice of using univariate robust ARIMA models (Chen and Liu, 1993) as in the HU model to forecast each coeficient $h$-steps ahead, $\hat{\beta}_{n,k,h}$, for each $k=1,\dots,K$. Subsequently, the forecasted log mortality rates can be calculated with:
@@
where $\hat{\sigma}^{2}_{\mu}(x)$ is the variance obtained from estimating the mean function, $u_{n+h|n,k}$ is the variance from the time series used to forecast $\hat{\beta}_{n+h|n,k}$, $v(x)$ are model fit errors which is estimated by averaging $\hat{e}_{t}(x)$, and finally $\sigma^{2}_{n+h|n}(x)$, the observational variance.
@@ 2.6.3 Bootstrap prediction interval
As the normality assumption is not satisfied (see A.1 in the Appendix) for the HUts model, we employ Hyndman and Shang (2009)’s method to construct the bootstrap prediction intervals.
The procedure begins by utilizing univariate time series models to generate multi-step-ahead forecasts for the scores of principal components, $\{\beta_{1,k}\}_{t=1}^{n}$. Subsequently, for $t=h+1,\dots,n$, we calculate the forecast errors for $h$-step-ahead forecasts as:
@@
The model’s residual, $\{e_{t}(x)\}$ is bootstrapped from the residuals $\{\hat{e}_{1}(x),\ldots,\hat{e}_{n}(x)\}$ by sampling with replacement. Similarly, smoothing errors $\hat{\epsilon}_{n+h,i}^{(l)}$ can be bootstrapped by resampling with replacement from $\{\hat{\epsilon}_{1,i},\ldots,\hat{\epsilon}_{n,i}\}$.
@@
$\hat{y}_{n+h|n}^{(l)}(x)=\hat{\mu}(x)+\sum_{k=1}^{K}\hat{\beta}_{n+h|n,k}^{(l)}Z(x)+\hat{e}_{n}^{(l)}+\hat{\sigma}_{n+h}(x)\hat{\epsilon}_{n+h,i}^{(l)}.$
The prediction intervals are then derived from these bootstrap variants using their percentiles.
Bias corrected bootstrap prediction intervals
The bootstrap method, while powerful, can introduce systematic biases due to its reliance on resampling operations. These biases may result from the resampling process not perfectly replicating the underlying population characteristics. Therefore, Efron and Tibshirani (1993) introduced the bias correction, to address these inaccuracies by adjusting the bootstrap intervals. This adjustment ensures that the intervals are correctly centered around the point forecasts, potentially producing more reliable intervals, particularly when the data does not conform to normal distribution assumptions (Efron and Tibshirani, 1993). This correction is applied to the bootstrap prediction intervals by a correction factor, $\hat{z}_{0}$, using the point forecasts.
$\hat{z}_{0}=\Phi^{-1}\left(\frac{\sum_{l=1}^{L}\mathbf{1}(\hat{y}_{n+h|n}^{(l)}(x)<\hat{y}_{n+h|n}(x))}{L}\right),$
where $\mathbf{1}(\cdot)$ is the indicator function, and $\Phi$ is the distribution function of the standard normal distribution for $l=1,\ldots,L$. This factor calculates the proportion of bootstrap variants that are less than the point forecast and then subsequently transformed using the inverse standard normal distribution function. Then, to determine the adjusted percentiles, we use:
$\alpha_{1}=\Phi(\hat{z}_{0}+z_{\alpha/2}),\quad\alpha_{2}=\Phi(\hat{z}_{0}+z_{1-\alpha/2}).$
These calculations yield the adjusted percentiles $\alpha_{1}$ and $\alpha_{2}$, which are used to construct the bias corrected prediction interval. This interval can effectively adjust for bias observed in the bootstrap sample, improving the accuracy of the predictive model’s interval estimates, and more accurately reflect the true variability and bias of the forecast.
@@ 3.2 Point forecast
The log mortality rates of 12 countries were forecasted with the HUts model using a truncation order $m=2$ (see A.2 in the Appendix). It is therefore intuitive to compare the forecast accuracy of the proposed method with the HU model, HUrob model, wHU model, and the LC model, which are commonly used on mortality rates of a single population. For the HUts model and the HU model variants, the number of principal components, $K$ is set to 6 which should be more than the required number of components considering the HU model is insensitive to the choice of $K$, presuming that $K$ is large enough (Hyndman and Booth, 2008). The models were fitted using an expanding window approach, incorporating data from the respective commencing years of each country as outlined in Table 1. The initial forecasting period covers the most recent two decades which ends at 2015. We fit this data using both models and produce forecasts up to ten-years ahead. The accuracy of these forecasts are then evaluated. Subsequently, the fitting period is then extended by one year and recalculate the forecasts for the same horizons. This procedure of expanding the training window continues until it includes data up to the year 2014.
In order to gauge the accuracy of the point forecast of the proposed signature mortality model, the mean squared error (MSE) and mean absolute error (MAE) are calculated and compared with the forecasts from the HU model. MSE quantifies the average squared difference between predicted and observed values, providing a measure of the overall accuracy and precision of predictions. On the other hand, MAE expresses the average difference between predicted and observed values. Both error measures can provide valuable insights into the model’s overall accuracy, with lower values indicating better forecasting performance. The MSE and MAE are computed using:
$MSE(h)$ $=\frac{1}{pq}\sum^{q}_{t=1}\sum^{p}_{i=1}(y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i})),$
$MAE(h)$ $=\frac{1}{pq}\sum^{q}_{t=1}\sum^{p}_{i=1}\lvert y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i})\rvert,$
@@
HUts HU HUrob wHU LC
Australia 0.00982 0.01027 0.01696 0.00952 0.10485
Belgium 0.02183 0.02144 0.02419 0.02047 0.18292
Bulgaria 0.01942 0.02131 0.02158 0.01968 1.08953
Denmark 0.03636 0.03810 0.04098 0.03590 0.11113
Finland 0.03856 0.04271 0.04435 0.03705 0.62800
France 0.00398 0.00642 0.00710 0.00399 0.15394
Ireland 0.05062 0.05174 0.05735 0.05161 0.10004
Italy 0.00711 0.00920 0.01600 0.00680 0.25442
Japan 0.00674 0.00747 0.00938 0.00687 0.11939
Netherlands 0.01242 0.01344 0.01569 0.01143 0.08333
Norway 0.04397 0.04303 0.04800 0.04264 0.56083
United States 0.00133 0.00207 0.00535 0.00129 0.01261
Mean 0.02101 0.02227 0.02558 0.02061 0.28341
00footnotetext: Note: Bold values indicate the smallest value in each row for the corresponding columns.
HUts HU HUrob wHU LC
Australia 0.06111 0.06614 0.09498 0.06036 0.23231
Belgium 0.08404 0.09092 0.10116 0.08193 0.30475
Bulgaria 0.09106 0.10052 0.09727 0.09138 0.57395
Denmark 0.11169 0.12029 0.12909 0.11127 0.22268
Finland 0.11279 0.12608 0.13552 0.10976 0.55369
France 0.04246 0.06100 0.06336 0.04256 0.28411
Ireland 0.12512 0.13419 0.14838 0.12719 0.22055
Italy 0.05330 0.06842 0.08936 0.05210 0.34423
Japan 0.05061 0.05701 0.06955 0.04975 0.27212
Netherlands 0.06802 0.07839 0.08692 0.06657 0.20192
Norway 0.11718 0.11890 0.13299 0.11479 0.52488
United States 0.02648 0.03459 0.05380 0.02628 0.08755
Mean 0.07866 0.08804 0.10020 0.07783 0.31856
00footnotetext: Note: Bold values indicate the smallest value in each row for the corresponding columns.
@@
Table 2 and Table 3 present the MSE and MAE for one-step-ahead forecasts for each country using the different models. The results indicate that while the wHU model often achieves the lowest MSE and MAE values for one-step-ahead forecasts, the HUts model still performs competitively, frequently achieving values close to the best-performing models. In comparison, the LC model shows significantly higher error rates, followed by the HU model and the HUrob model.
HUts HU HUrob wHU LC
Australia 0.01835 0.02245 0.02890 0.01642 0.11772
Belgium 0.02627 0.02863 0.03211 0.02620 0.20303
Bulgaria 0.04595 0.05212 0.04574 0.05478 0.59795
Denmark 0.04891 0.05484 0.05764 0.05007 0.08655
Finland 0.04698 0.05988 0.05328 0.04538 0.69713
France 0.00882 0.02290 0.01489 0.00978 0.15429
Ireland 0.07035 0.07198 0.07707 0.07785 0.11353
Italy 0.01625 0.03153 0.02982 0.01546 0.25481
Japan 0.01045 0.01288 0.01361 0.01020 0.21069
Netherlands 0.01785 0.02171 0.02374 0.01859 0.07366
Norway 0.05056 0.05309 0.05856 0.05053 0.58555
United States 0.00746 0.00900 0.01252 0.00896 0.01848
Mean 0.03068 0.03675 0.03732 0.03202 0.25945
00footnotetext: Note: Bold values indicate the smallest value in each row for the corresponding columns.
HUts HU HUrob wHU LC
Australia 0.09131 0.11006 0.12925 0.08682 0.25691
Belgium 0.10163 0.11696 0.12482 0.10307 0.32362
Bulgaria 0.15511 0.17190 0.15103 0.15636 0.44353
Denmark 0.14989 0.16413 0.17112 0.15113 0.20720
Finland 0.14040 0.17654 0.15844 0.13800 0.59051
France 0.07102 0.12234 0.09527 0.07427 0.29203
Ireland 0.16031 0.18212 0.18869 0.17752 0.24991
Italy 0.09528 0.14200 0.13301 0.08831 0.35423
Japan 0.07659 0.08745 0.08784 0.07277 0.36327
Netherlands 0.09511 0.11317 0.11668 0.09702 0.19762
Norway 0.13801 0.15506 0.16166 0.13688 0.54332
United States 0.06428 0.07333 0.08508 0.06843 0.10677
Mean 0.11158 0.13459 0.13357 0.11255 0.32741
00footnotetext: Note: Bold values indicate the smallest value in each row for the corresponding columns.
@@
For five-step-ahead forecasts, the HUts model demonstrates more competitive performance. Table 4 and Table 5 show that overall the HUts model achieves lower error metrics compared to other models, particularly in contrast to the LC model. This suggests an improvement in the HUts model’s forecasting ability as the horizon increases.
HUts HU HUrob wHU LC
Australia 0.0418 0.0511 0.0509 0.0351 0.1432
Belgium 0.0364 0.0380 0.0450 0.0404 0.2171
Bulgaria 0.0976 0.0931 0.0811 0.1280 0.1733
Denmark 0.0764 0.0825 0.0927 0.0825 0.0767
Finland 0.0653 0.0894 0.0773 0.0617 0.7650
France 0.0201 0.0526 0.0278 0.0242 0.1474
Ireland 0.1111 0.1297 0.1177 0.1413 0.1320
Italy 0.0403 0.0603 0.0594 0.0433 0.2481
Japan 0.0348 0.0399 0.0289 0.0362 0.3664
Netherlands 0.0326 0.0391 0.0406 0.0377 0.0513
Norway 0.0695 0.0775 0.0802 0.0718 0.5473
United States 0.0226 0.0196 0.0342 0.0305 0.0247
Mean 0.0540 0.0644 0.0613 0.0611 0.2410
00footnotetext: Note: Bold values indicate the smallest value in each row for the corresponding columns.
HUts HU HUrob wHU LC
Australia 0.1483 0.1705 0.1749 0.1383 0.2937
Belgium 0.1323 0.1421 0.1558 0.1434 0.3430
Bulgaria 0.2490 0.2389 0.2093 0.2341 0.2801
Denmark 0.2097 0.2182 0.2352 0.2151 0.2077
Finland 0.1804 0.2334 0.2105 0.1754 0.6260
France 0.1150 0.1905 0.1360 0.1204 0.2950
Ireland 0.2164 0.2663 0.2510 0.2641 0.2905
Italy 0.1623 0.2056 0.2019 0.1550 0.3655
Japan 0.1410 0.1536 0.1289 0.1435 0.4753
Netherlands 0.1448 0.1601 0.1637 0.1564 0.1820
Norway 0.1809 0.2089 0.2087 0.1837 0.5392
United States 0.1121 0.1122 0.1380 0.1250 0.1259
Mean 0.1660 0.1917 0.1845 0.1712 0.3353
00footnotetext: Note: Bold values indicate the smallest value in each row for the corresponding columns.
@@
The results for ten-step-ahead forecasts, shown in Table 6 and Table 7, further confirm the strong performance of the HUts model. It consistently achieves lower or comparable error metrics, reinforcing its effectiveness in long-term mortality rate forecasting. In most countries, the HUts model has lower MSE and MAE values, with exceptions in a few cases such as the United States, where the HU model occasionally performs better. This consistent performance across a majority of countries highlights the robustness of the HUts model in capturing long-term trends in mortality rates.
@@
Across different forecasting horizons, the best model is generally a choice between the HUts and wHU models, as these models tend to have the lowest error metrics. The wHU model performs better at shorter forecast horizons, as seen in the one-step-ahead results. However, as the forecasting horizon increases, the HUts model demonstrates improved performance and eventually surpasses the wHU model. This shift indicates that the HUts model’s accuracy and reliability become more pronounced with longer-term forecasts.
@@
Notably, in the cases of France and Ireland, the HUts model performs consistently well across all forecasting horizons, maintaining lower MSE and MAE values compared to other models. Conversely, for Australia, the HUts model does not appear to be the most favorable at any point across the forecast horizons, with other models, such as wHU, often yielding better results. The following section will take a closer look at the mortality rate forecasts for France and Australia, exploring the nuances and potential reasons behind the differing performances of the HUts model against the HU variants. The LC model is excluded to ensure the plot scales remain appropriate and comparisons between the other models are more visible.
@@ 3.2.1 A closer look at the French and Australian mortality data
Figure 3: (a) MSE and (b) MAE plots for French mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models
Figure 4: (a) MSE and (b) MAE plots for Australian mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models
Figure 5: One-step-ahead (a) MSE and (b) MAE (c) ME plots averaged over years in forecasting period of French mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models. A horizontal line at 0 is included in the ME plot to facilitate the assessment of forecast bias
Figure 6: Five-step-ahead (a) MSE and (b) MAE (c) ME plots averaged over years in forecasting period of French mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models. A horizontal line at 0 is included in the ME plot to facilitate the assessment of forecast bias
Figure 7: Ten-step-ahead (a) MSE and (b) MAE (c) ME plots averaged over years in forecasting period of French mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models. A horizontal line at 0 is included in the ME plot to facilitate the assessment of forecast bias
Figure 8: One-step-ahead (a) MSE and (b) MAE (c) ME plots averaged over years in forecasting period of Australian mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models. A horizontal line at 0 is included in the ME plot to facilitate the assessment of forecast bias
Figure 9: Five-step-ahead (a) MSE and (b) MAE (c) ME plots averaged over years in forecasting period of Australian mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models. A horizontal line at 0 is included in the ME plot to facilitate the assessment of forecast bias
Figure 10: Ten-step-ahead (a) MSE and (b) MAE (c) ME plots averaged over years in forecasting period of Australian mortality of the HUts (red solid line), HU (blue dashed line), HUrob (green dotted line), and the wHU (orange dash-dotted line) models. A horizontal line at 0 is included in the ME plot to facilitate the assessment of forecast bias
To assess the robustness of the HUts model, we examine the mortality rates of France and Australia more closely. Figures 1 (a) and 2 (a) include the observed log mortality rates, and Figures 1 (b) and 2 (b) show the smoothed log mortality rates with red representing the earlier years. Figures 1 (c) and 2 (c) the smoothed log mortality rates viewed as a time series with red representing the earlier ages. France’s mortality data, notably influenced by periods of war and disease outbreaks, presents distinct outliers which are evident in the volatility observed in their mortality trends, as shown in Figure 1 which are most visible around the 1920s and 1940s. This provides an oppurtunity to test for the HUts model’s ability to handle anomalies in data. In contrast, Australia’s mortality data, starting from 1921 appears to have a smoother trend, without any significant outliers as depicted in Figure 2.
@@
The ability of the HUts model to effectively handle anomalies in French mortality is illustrated in Figure 3, where the model demonstrates a substantial reduction in MSE and MAE across various forecast horizons when compared to the HU model. This significant improvement shows the HUts model’s robustness against data irregularities, making it a reliable choice for forecasting in scenarios with potential outliers. Additionally, the HUts model showcases lower MSE and MAE than both the HUrob and wHU models, further affirming its superior performance in handling French mortality data. Conversely, Australia presents a more stable dataset with mortality rates following a consistent trend and lacking significant outliers. The MSE and MAE plots for Australia in Figure 4 show less pronounced differences compared to France. While the HUts model demonstrates superior performance over the HU and HUrob models, affirming its effectiveness even in more uniform datasets, the wHU model performs better than the HUts model. This suggests that the wHU model might be more suited for datasets with consistent trends and fewer irregularities. This comparative analysis highlights the importance of selecting appropriate models based on the characteristics of the dataset, with the HUts model excelling in more variable datasets and the wHU model in more stable ones. The MSE and MAE plots of the remaining 10 countries can be found at A.3 in the Appendix.
@@
To gain deeper insights into the model performance, the MSE and MAE values are considered across the age spectrum. Instead of using Equation (3.2), the MSE and MAE values are averaged only over the years of the forecasting period. This approach will allow us to observe how the models perform at different age groups, providing a more granular understanding of the forecast accuracy. Additionally, we will also calculate the mean error (ME) to identify whether the models tend to overpredict or underpredict mortality rates at specific ages. The MSE, MAE and ME are computed with the following:
$MSE(h,x_{i})$ $=\frac{1}{q}\sum_{t=1}^{q}(y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i}))^{2},$
$MAE(h,x_{i})$ $=\frac{1}{q}\sum_{t=1}^{q}\lvert y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i})\rvert,$
$ME(h,x_{i})$ $=\frac{1}{q}\sum_{t=1}^{q}(y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i})).$
@@
Upon a closer inspection of the evaluation metrics across different age groups of the French mortality data in Figures 5, 6, and 7, the HUts model generally surpasses every other model in terms of forecasting performance. For one-step-ahead forecasts (Figure 5), the HUts and wHU models show similar performance in terms of MSE and MAE, with both models outperforming the HU and HUrob models. The HUts model particularly excels in forecasting ages between 20 and 40, where it demonstrates substantially lower error values compared to the wHU model. The ME plot indicates that the HUts model’s forecasts are mostly unbiased across ages, though there are some dips around age 15 and peaks around ages 20 and 100. The wHU model shows a similar trend, while the HU and HUrob models generally underpredict mortality. For five-step-ahead forecasts (Figure 6), the HUts model continues to perform well, maintaining lower MSE and MAE values across most age groups compared to the other models. The HUts model shows improved performance at middle ages (30-50), while the HU and HUrob models, although better at these ages, struggle during early and older ages, contributing to their higher overall error metrics. The ME plot for the HUts model remains close to zero, indicating minimal bias, whereas the wHU model exhibits similar trends, and the HU and HUrob models continue to underpredict mortality. For ten-step-ahead forecasts (Figure 7), the HUts model’s performance remains robust, achieving lower MSE and MAE values across most age groups. However, it still faces competition from the wHU model, which performs better in certain age ranges. The ME plot for the HUts model shows a slight bias at extreme ages, but overall, it maintains a near-zero mean error across most ages, suggesting accurate and unbiased forecasts. The HU and HUrob models exhibit larger biases, contributing to their higher error metrics. Overall, these observations suggest that while the HUts model performs exceptionally well, particularly as the forecast horizon increases, the wHU model also shows strong performance, especially at shorter forecast horizons. The choice between these models often depends on the specific age group and forecasting horizon, with the HUts model becoming more favorable as the horizon increases.
@@
Similarly, the evaluation metrics for the Australian mortality data in Figures 8, 9, and 10 highlight the competitive nature of the HUts model, particularly at shorter forecast horizons. For one-step-ahead forecasts (Figure 8), the HUts model performs comparably to the wHU model, with both showing nearly identical MSE and MAE values. This indicates that the HUts model is highly effective at short-term forecasting. The ME plot reveals minimal bias across ages for the HUts model, similar to the wHU model, while the HU and HUrob models exhibit a tendency to underpredict mortality. At the five-step-ahead horizon (Figure 9), the HUts model continues to perform well but begins to show higher MSE and MAE values in the 20-40 age range compared to the wHU model. This underperformance in mid-age groups becomes more evident, reflecting some biases that start to emerge at this forecast horizon. The ME plot for the HUts model indicates underprediction biases in this age range, contributing to the increased error metrics. By the ten-step-ahead forecasts (Figure 10), the HUts model, although still competitive, shows further increased MSE and MAE values around ages 20-40. This trend underscores the model’s underprediction in this specific age group, significantly affecting the overall error metrics. Despite this, the HUts model maintains reasonable performance across other age groups, yet the wHU model generally achieves better results at these longer horizons. While the HUts model shows strong performance and remains competitive across various forecasting horizons, it tends to exhibit higher MSE and MAE values around ages 20-40, especially at longer horizons, due to underprediction. This suggests that the wHU model might be better suited for datasets with more consistent trends and fewer irregularities, particularly for longer-term forecasts.

### v2

@@ 2.6.1 Point forecast
Similar to that of the HU model, the time dependent coefficients are extrapolated by fitting them into univariate time series $\{\hat{\beta}_{t,k}\}_{t=1}^{n}$. We preserve the choice of using univariate ARIMA models as in the HU model to forecast each coefficient $h$-steps-ahead, $\hat{\beta}_{n+h|n,k}$, for each $k=1,\dots,K$. Subsequently, the forecasted log mortality rates can be calculated with:
@@
where $\hat{\sigma}^{2}_{\mu}(x)$ is the variance obtained from estimating the mean function, $u_{n+h|n,k}$ is the variance from the time series used to forecast $\hat{\beta}_{n+h|n,k}$, $v(x)$ are model fit errors estimated by averaging $\hat{e}_{t}(x)$, and finally $\sigma^{2}_{n+h|n}(x)$, the observational variance.
@@ 2.6.3 Bootstrap prediction interval
Since the HUts model, like the wHU model, do not satisfy the normality assumption (see A.1 in the Appendix), we employ the same method used to construct bootstrap prediction intervals for the wHU model in Hyndman and Shang (2009) to construct prediction intervals for the HUts model.
The procedure begins by utilizing univariate time series models to generate multi-step-ahead forecasts for the scores of principal components, $\{\beta_{t,k}\}_{t=1}^{n}$. Subsequently, for $t=h+1,\dots,n$, we calculate the forecast errors for $h$-step-ahead forecasts as:
@@
The model’s residual, $\hat{e}_{n+h|n}^{(l)}(x)$ is bootstrapped from the residuals $\{\hat{e}_{1}(x),\ldots,\hat{e}_{n}(x)\}$ by sampling with replacement. Similarly, smoothing errors $\hat{\epsilon}_{n+h,i}^{(l)}$ can be bootstrapped by resampling with replacement from $\{\hat{\epsilon}_{1,i},\ldots,\hat{\epsilon}_{n,i}\}$.
@@
$\hat{y}_{n+h|n}^{(l)}(x_{i})=\hat{\mu}(x)+\sum_{k=1}^{K}\hat{\beta}_{n+h|n,k}^{(l)}\hat{Z}_{k}(x_{i})+\hat{e}_{n+h|n}^{(l)}(x_{i})+\hat{\sigma}_{n+h}(x_{i})\hat{\epsilon}_{n+h,i}^{(l)}.$
Subsequently, the $(1-\alpha)100\%$ prediction intervals, $[l_{h}(x),u_{h}(x)]$ can then be obtained from the $(1-\alpha/2)$ and $(\alpha/2)$ quantiles of these bootstrap variants. We also consider the adjustment made to the bootstrap prediction intervals in Hyndman and Shang (2009):
$[0.5\{l_{h}(x)+u_{h}(x)\}-\{u_{h}(x)-l_{h}(x)\}p(x),0.5\{l_{h}(x)+u_{h}(x)\}+\{u_{h}(x)-l_{h}(x)\}p(x)]$
to account for the uncertainty in our forecast by calibrating the intervals using one-step-ahead forecast errors. Here, $p(x)=d(x)/[u_{1}(x)-l_{1}(x)]$, where $d(x)$ is the empirical differences between the $(1-\alpha/2)$ and $(\alpha/2)$ quantiles of the in-sample one-step-ahead forecast errors $\{\hat{f}_{t+1}(x)-\hat{f}_{t+1|t}(x);t=r,\cdots,n-1\}$, and $r$ is the smallest number of observations used to fit the model.
@@ 3.2 Point forecast
The log mortality rates of 12 countries were forecasted with the HUts model using a truncation order $m=2$ (see A.2 in the Appendix). It is therefore intuitive to compare the forecast accuracy of the proposed method with the HU model, HUrob model, and the wHU model, which are commonly used on mortality rates of a single population. The selection of the number of principal components, $K$ can impact the forecasting performance of the HUts model and the HU model variants. They can be chosen based off a threshold function that measures the proportion of variance explained by each principal component, or by minimizing the integrated squared forecast error (Hyndman and Ullah, 2007). However, Hyndman and Booth (2008) noted that the HU models are relatively insensitive to the choice of $K$ as long as it is sufficiently large, and that setting $K=6$ is more than adequate to capture the essential features of the data. For consistency across all models, we adopt $K=6$ for both the HU model variants and the proposed HUts model. The models were fitted using an expanding window approach, incorporating data from the respective commencing years of each country as outlined in Table 1. The initial forecasting period covers the most recent two decades which ends at 2015. We fit this data using both models and produce forecasts up to ten years ahead. The accuracy of these forecasts are then evaluated. Subsequently, the fitting period is then extended by a year and the forecasts for the same horizons are recalculated. This procedure of expanding the training window continues until it includes data up to the year 2014.
In order to gauge the accuracy of the point forecast of the HUts model, the mean squared errors (MSE) are calculated and compared with those of the forecasts of the HU model variants. MSE quantifies the average squared difference between the predicted and observed values, providing a measure of the overall accuracy and precision of predictions with lower values indicating better forecasting performance. The MSE is computed using:
$MSE(h)=\frac{1}{pq}\sum^{q}_{t=1}\sum^{p}_{i=1}(y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i}))^{2},$ (3.1)
@@
Figure 1: Root MSE of (a) one, (b) five, and (c) ten-step-ahead point forecasts of log mortality rates by model and country
@@
Figure 1 presents the plot of the root MSE of $h=1,5,10$ forecasts across the different models. The square roots of the MSE values are plotted instead of the MSE themselves to ensure the plot scale accommodates all data. Readers are directed to A.4 in the Appendix for the tabulated MSE values of all countries. The results indicate that while the wHU model often achieves the lowest MSE values for one-step-ahead forecasts, the HUts model still performs competitively, frequently achieving values close to the best-performing models. In comparison, the HU model shows higher error rates, followed by the HUrob model.
@@
For five-step-ahead forecasts, the HUts model demonstrates more competitive performance. Figure 1 (b) shows that overall the HUts model achieves lower error metrics compared to the other models, suggesting an improvement in the HUts model’s forecasting ability as the forecast horizon increases.
@@
The results for ten-step-ahead forecasts, shown in Figure 1 (c), further confirms the strong performance of the HUts model. It consistently achieves lower or comparable MSE, reinforcing its effectiveness in long-term mortality rate forecasting. In most countries, the HUts model has lower MSE values, with exceptions to a few cases such as in the United States, where the HU model occasionally performs better. This consistent performance across a majority of countries highlights the robustness of the HUts model in capturing long-term trends in mortality rates.
@@
Across different forecasting horizons, the best model is generally a choice between the HUts and wHU models, as these models tend to have the lowest MSE values. The wHU model performs better at shorter forecast horizons, as seen in Figure 1 (a). However, as the forecasting horizon increases, the HUts model demonstrates improved performance and eventually surpasses the wHU model. This shift indicates that the HUts model’s accuracy and reliability become more pronounced with longer-term forecasts.
@@
Notably, in the cases of France and Ireland, the HUts model performs consistently well across all forecasting horizons, maintaining lower MSE values compared to other models. Conversely, for Australia, the HUts model does not appear to be the most favorable at any point across the forecast horizons, with other models, such as the wHU model, often yielding better results. The following section will take a closer look at the mortality rate forecasts for France and Australia, exploring the nuances and potential reasons behind the differing performances of the HUts model against the HU model variants.
@@ 3.2.1 A closer look at the French and Australian mortality data
To assess the robustness of the HUts model, we examine the mortality rates of France and Australia more closely. Figures 2 (a) and 3 (a) include the observed log mortality rates, and Figures 2 (b) and 3 (b) show the smoothed log mortality rates with red representing the earlier years. Figures 2 (c) and 3 (c) the smoothed log mortality rates viewed as a time series with red representing the earlier ages. France’s mortality data, notably influenced by periods of war and disease outbreaks, presents distinct outliers which are evident in the volatility observed in their mortality trends, as shown in Figure 2 which are most visible around the 1920s and 1940s. This provides an oppurtunity to test for the HUts model’s ability to handle anomalies in data. In contrast, Australia’s mortality data, starting from 1921 appears to have a smoother trend, without any significant outliers as depicted in Figure 3.
@@
Figure 4: MSE plots for (a) French and (b) Australian mortality of the HUts, wHU, HUrob, and HU models
@@
The ability of the HUts model to effectively handle anomalies in French mortality is illustrated in Figure 4 (a), where the model demonstrates a substantial reduction in MSE across various forecast horizons when compared to the HU model. This significant improvement shows the HUts model’s robustness against data irregularities, making it a reliable choice for forecasting in scenarios with potential outliers. Additionally, the HUts model showcases lower MSE than both the HUrob and wHU models, further affirming its superior performance in handling French mortality. Conversely, Australia presents a more stable dataset with mortality rates following a consistent trend and lacking significant outliers. The MSE plot for Australia in Figure 4 (b) shows less pronounced differences compared to France. While the HUts model demonstrates superior performance over the HU and HUrob models, affirming its effectiveness even in more uniform datasets, the wHU model performs better than the HUts model. This suggests that the wHU model might be more suited for datasets with consistent trends and fewer irregularities. This comparative analysis highlights the importance of selecting appropriate models based on the characteristics of the dataset, with the HUts model excelling in more variable datasets and the wHU model in more stable ones. The MSE plots of the remaining countries can be found in A.5 in the Appendix.
@@
To gain deeper insights into the model performance, the MSE values are considered across the age spectrum. Instead of using Equation (3.1), the MSE values are averaged only over the years of the forecasting period. This approach will allow us to observe how the models perform at different age groups, providing a more granular understanding of the forecast accuracy. The MSE across ages are computed with the following:
$MSE(h,x_{i})=\frac{1}{q}\sum_{t=1}^{q}(y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i}))^{2}.$
Figure 5: One-step-ahead forecast MSE averaged over years in forecasting period of (a) French and (b) Australian mortality for the HUts, wHU, HUrob, and HU models
Figure 6: Five-step-ahead forecast MSE averaged over years in forecasting period of (a) French and (b) Australian mortality for the HUts, wHU, HUrob, and HU models
Figure 7: Ten-step-ahead forecast MSE averaged over years in forecasting period of (a) French and (b) Australian mortality for the HUts, wHU, HUrob, and HU models
@@
Upon a closer inspection of the MSE across different age groups of the French mortality data in Figures 5 (a), 6 (a), and 7 (a), the HUts model generally surpasses every other model in terms of forecasting performance. For one-step-ahead forecasts (Figure 5 (a)), the HUts and wHU models show similar performance in terms of MSE, with both models outperforming the HU and HUrob models substantially. For five-step-ahead forecasts (Figure 6 (a)), the HUts model continues to perform well, maintaining lower MSE values across most age groups compared to the other models. The HUts model shows improved performance at middle ages (30-50) over the wHU model, while the HU and HUrob models, although better at these ages, struggle during early and older ages, contributing to their higher overall error metric. For ten-step-ahead forecasts (Figure 7 (a)), the HUts model’s performance remains robust, achieving lower MSE values across most age groups. However, it still faces competition from the wHU model, which performs better in certain age ranges. Overall, these observations suggest that while the HUts model performs exceptionally well, particularly as the forecast horizon increases, the wHU model also shows strong performance, especially at shorter forecast horizons. The choice between these models often depends on the specific age group and forecasting horizon, with the HUts model becoming more favorable as the horizon increases.
@@
Similarly, MSE for the Australian mortality forecasts in Figures 5 (b), 6 (b), and 7 (b) highlight the competitive nature of the HUts model, particularly at shorter forecast horizons. For one-step-ahead forecasts (Figure 5 (b)), the HUts model performs comparably to the wHU model, with both showing nearly identical MSE values. This indicates that the HUts model is effective at short-term forecasting. At the five-step-ahead horizon (Figure 6 (b)), the HUts model continues to perform well but begins to show higher MSE values in the 20-40 age range compared to the wHU model. This underperformance in mid-age groups becomes more evident, reflecting some potential biases that start to emerge at this forecast horizon. By the ten-step-ahead forecasts (Figure 7 (b)), the HUts model, although still competitive, shows further increased MSE values around ages 20-40. This trend highlights the model’s underestimation (see A.6 of the Appendix) in this specific age group, significantly affecting the overall error metric. Despite this, the HUts model maintains reasonable performance across other age groups, yet the wHU model generally achieves better results at these longer horizons, suggesting that the wHU model might be better suited for datasets with more consistent trends and fewer irregularities.